

********************************************************************
UARS WEEKLY STATUS REPORT - 25 FEB 2005
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
59% of each day's data amounting to about 507 minutes of solar
observing time per day.
***********
SUSIM
***********
During this past week, SUSIM successfully gathered daily mid (1.1 nm)
resolution, MgII, occultation, solar pointing offset, and continuous
monitoring scans for time periods with UARS telemetry or tape
recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004289 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004290 and 2004299 are currently waiting archival at GSFC DAAC.
HEPSA data between 2004300 and 2004309 are currently being processed.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004299. MEPS data between 2004300 and 2004309
are currently waiting archival at GSFC DAAC.
PEM continues to receive NMC data files from NOAA. NMC host machine
changed during mid-January which caused some NMC data transfer failures;
however, this situation was corrected and recovery completed with the
assistance of Roger Lin at NOAA. No problems have been encountered since
updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 February 2005 (Orbit 73466, GMT Day 042) through
18 February 2005 (Orbit 73571, GMT Day 049).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 045/1254 73504 DAY ONLY TO FULLTIME
SUSIM 045/1616 73506 POWER ON
PEM 047/1248 73534 POWER ON - DAY ONLY OPS
HALOE 048/1246 73549 POWER ON
A Forward-to-Reverse Yaw Around maneuver (number 137) was successfully
executed in Orbit 73476 beginning 043/ 15:19:00Z and completing at
043/ 15:56:10Z on 12 February 2005. All instruments were in their off
or standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 56.85 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.2 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 19.8 and 18.8 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 21
March 2005 (DOY 080).
The beta angle ranged between the angles of 0.9 degrees to 22.5
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 63.0 degrees on 02 March 2005 (DOY 061).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 3.27 to 3.62 degrees C. Additional
battery performance data for each battery on 11 February 2005 and 18
February 2005 are provided in Appendix D.
The clock error ranged from 10.3 to -10.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
SOLSTICE = 01
SUSIM = 01
PEM = 01
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to monitor the batteries closely. Over
the last few weeks a slow increasing trend has been seen in the
battery temperatures and the differential voltage on Battery-2. Then
on 17 February 2005, after the HALOE turn on, Battery-3 differential
voltage increased dramatically. This was alarming at first but after
4 orbits the differential voltage came back into nominal range. Since
the batteries are still requiring more charge, the MPS continues to
utilize the VT-7 charge rate.
The FOT has completed their CMS software acceptance testing. All
tests results were positive and the official release of the new
software was installed on the backup CMS. With the install complete,
the FOT will run a few simulated days of operations to confirm the
software. If all goes well, then the new release will be put onto the
operational CMS.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 March 2005 (DOY 080). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 February 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73475 043/1401 05-042 MACS YAW:TAM TO MONITOR
73476 043/1508 05-042 MACS YAW:REV ALIGN/DIRECTION CHANGE
73476 043/1623 05-042 MACS YAW:TAM TO UPDATE
73495 044/2236 05-043 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73504 045/1254 05-044 SOLSTICE DAY ONLY TO FULLTIME OPS
73506 045/1616 05-045 SUSIM POWER ON - PART ONE
73517 046/0928 05-046 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
73519 046/1251 05-045 SUSIM POWER ON - PART TWO
73534 047/1248 05-047 PEM POWER ON - DAY ONLY OPS
73549 048/1246 05-048 HALOE POWER ON - PART ONE
73551 048/1606 05-049 HALOE POWER ON - PART TWO
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 February 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
OBSPMACOOL23 PMON ACTION 23 73553
OBSPMACOOL31 PMON ACTION 31 73553
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 18 February 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73495 044/2242 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73517 046/0817 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 February 2005
11 February 2005 - GMT Day 042 Beta = 05.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 67.2/ 11.2 +8.40/+7.74 +28.0/-12.8 27.0 1.06 18.8
3 -33.6/-112.0 +5.79/+5.15 +25.6/-12.0 26.9 1.04 17.8
18 February 2005 - GMT Day 049 Beta = 22.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.11/-3.05 +00.0/+00.0 07.0 0.00 0.0
2 235.2/ 67.2 +10.06/+9.06 +28.8/-12.8 26.7 1.07 19.9
3 -16.8/-235.2 +6.44/+5.79 +25.6/-12.0 26.7 1.05 18.4
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 February 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT 73479 043/2107 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-008 FOT 73431 040/1517 38482 IP NOC Router Problem
(:37)
05-009 FOT 73448 041/1836 N/A Mutual Interference
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT See List Feb 05 30832 Generic Late Acquisition
05-002 MPS 73008 012/1251 N/A BATT 3 Diff Voltage High